# Visual Planning: Let’s Think Only with Images

### v1

@@ Abstract
Recent advancements in Large Language Models (LLMs) and their multimodal extensions (MLLMs) have substantially enhanced machine reasoning across diverse tasks. However, these models predominantly rely on pure text as the medium for both expressing and structuring reasoning, even when visual information is present. In this work, we argue that language may not always be the most natural or effective modality for reasoning, particularly in tasks involving spatial and geometrical information. Motivated by this, we propose a new paradigm, Visual Planning, which enables planning through purely visual representations, independent of text. In this paradigm, planning is executed via sequences of images that encode step-by-step inference in the visual domain, akin to how humans sketch or visualize future actions. We introduce a novel reinforcement learning framework, Visual Planning via Reinforcement Learning (VPRL), empowered by GRPO for post-training large vision models, leading to substantial improvements in planning in a selection of representative visual navigation tasks, FrozenLake, Maze, and MiniBehavior. Our visual planning paradigm outperforms all other planning variants that conduct reasoning in the text-only space. Our results establish Visual Planning as a viable and promising alternative to language-based reasoning, opening new avenues for tasks that benefit from intuitive, image-based inference. Code is available at: https://github.com/yix8/VisualPlanning.
@@ 1 Introduction
Building on this observation, we argue that performing multimodal reasoning only in the text pathway may not always offer the most intuitive or effective strategy, particularly for tasks that depend heavily on visual information and/or are ‘vision-first’ by design. Indeed, recent results from multimodal benchmarks [43, 30, 7, 8] offer growing evidence that purely language-based reasoning falls short in certain domains, particularly those involving spatial, geometric, or physical dynamics [56]. Such reliance on grounding visual information into text before reasoning introduces a modality gap that hinders the model’s ability to capture visual features and state transitions. This highlights a potential shortcoming of current MLLMs: while they process image inputs, they do not naturally “think” in images. For instance, tasks such as planning a route through a maze, designing the layout of a room, or predicting the next state of a mechanical system are often better served by visual representations, as verbal descriptions may struggle to accurately capture complex spatial reasoning relationships. These examples suggest a broader question, which we aim to tackle in this work: can models plan in non-verbal modalities, such as images, without being mediated by text?
@@
We validate the feasibility of our paradigms on grid-based navigation as a representative of spatial planning tasks, including Maze [23], FrozenLake [53], and MiniBehavior [25], where one agent is requested to navigate to a target location successfully without violating environment constraints. Our experiments reveal that the visual planning paradigm substantially surpasses the traditional textual reasoning method by supervised fine-tuning (SFT), achieving more than 40% higher average exact-match rate. In addition to better performance, our novel method VPRL exhibits stronger generalization to out-of-distribution scenarios than the SFT method in the visual planning paradigm (VPFT). To the best of our knowledge, we are the first to apply RL to image generation in the context of planning; the main contributions comprise the following:
@@ 2.1 The Visual Planning Paradigm
The majority of prior visual reasoning benchmarks [14, 1, 55] can be and is typically tackled by grounding the visual information in the textual domain [18, 40, 57], followed by a few steps of textual reasoning. However, once the visual content is mapped to text (e.g., object names, attributes, or relations), the problem gets reduced to a language reasoning task, where the reasoning is carried out by the language model, even without reflecting any information from the visual modality during the reasoning process.
Our visual planning paradigm is fundamentally different. It performs planning purely within the visual modality. We formally define visual planning as a process of generating a sequence of intermediate images $\mathcal{T}=(\hat{v}_{1},\ldots,\hat{v}_{n})$, where each $\hat{v}_{i}$ represents a visual state that together constitute a visual planing trajectory, given the input image $v_{0}$. Specifically, let $\pi_{\theta}$ denote a generative vision model parameterized by $\theta$. The visual planning trajectory $\mathcal{T}$ is generated autoregressively, where each intermediate visual state $\hat{v}_{i}$ is sampled conditioned on the initial state and previously generated states:
$\hat{v}_{i}\sim\pi_{\theta}(v_{i}|v_{0},\hat{v}_{1},...,\hat{v}_{i-1})$ (1)
@@ 2.2 Reinforcement Learning for Large Vision Models
Reinforcement learning (RL) has shown notable advantages in improving the generalization of autoregressive models by optimizing with sequence-level rewards beyond token-level supervision signals [11]. In autoregressive image generation, an image is represented as a sequence of visual tokens. Inspired by the success of RL in language reasoning [16], we introduce an RL-based training framework for visual planning empowered by Group Relative Policy Optimization (GRPO) [44]. It leverages the transitions between visual states to compute the reward signals while verifying the constraints from the environments. To enforce the policy model that generates valid actions with diverse exploration during the RL process, we then propose a novel two-stage reinforcement learning framework for visual planning. In Stage 1, we first apply supervised learning to initialize the policy model with random walks. Model’s visual planning is then optimized by the RL training in Stage 2.
Stage 1: Policy Initialization. In this stage, we initialize the model $\pi_{\theta}$ by training it on random trajectories obtained by random walks in the environment. The goal here is to generate valid sequences of visual states and retain exploration capability in a ‘simulated’ environment. For training, each trajectory consists of a sequence of visual states $(v_{0},\ldots,v_{n})$. From each trajectory, we extract $n-1$ image pairs of the form $(v_{\leq i},v_{i+1})$, where $v_{\leq i}$ represents the prefix sequence $(v_{0},\ldots,v_{i})$. Subsequently, given an input prefix, the model is exposed to a set of plausible next states $\{v_{i+1}^{(j)}\}_{j=1}^{K}$, collected from $K$ valid trajectories that share the same prefix. To prevent overfitting to the specific transition and encourage stochasticity, we randomly sample one candidate $v_{i+1}^{(\ell)}$ from this set at each training step as the supervision target by minimizing the following loss function of visual planning via fine-tuning (VPFT):
$\mathcal{L}_{\text{VPFT}}(\theta)=-\mathbb{E}_{(v_{\leq i},\,v_{i+1}^{(\ell)})}\Bigl{[}\log\pi_{\theta}\!\bigl{(}v^{(\ell)}_{i+1}\,\big{|}\,v_{\leq i}\bigr{)}\Bigr{]}.$ (2)
@@
Stage 2: Reinforcement Learning for Visual Planning. Building on Stage 1, where the model is initialized with random trajectories, it acquires the effective exploration capability. This property is essential for RL, as it ensures coverage over all possible transitions and prevents collapse to suboptimal behaviors. Stage 2 then leverages this ability to simulate the outcomes of potential actions by generating the next visual state and guiding the model to effectively do the planning. During this stage, the RL algorithm provides feedback and rewards based on the correctness of the simulated actions, gradually enabling the model to learn effective visual planning. Specifically, given an input prefix $v_{\leq i}$, the behavior model $\pi_{\theta_{\text{old}}}$ samples a group of $G$ candidate responses $\{\hat{v}_{i+1}^{(1)},\ldots,\hat{v}_{i+1}^{(G)}\}$. Each response represents a hypothetical visual state corresponding to a planned action $a_{i}^{(k)}$ at time step $i$. To interpret these transitions, we employ a rule-based parsing function that maps pairs of visual states $(v_{i},\hat{v}_{i+1}^{(k)})$ to discrete actions. The candidate response is then scored using a composite reward function $r(v_{i},\hat{v}_{i+1}^{(k)})$, which quantifies whether the generated visual state represents meaningful progress toward the goal state. The reward design is described in detail in the next paragraph.
Instead of relying on a learned critic to estimate value functions which may introduce additional sources of uncertainty and complexity, GRPO provides more computationally efficient and interpretable training signals by computing relative advantages through comparisons within the group. In this case, the relative advantage of each candidate is $A^{(k)}=\frac{r^{(k)}-\text{mean}\left\{r^{(1)},r^{(2)},\ldots,r^{(G)}\right\}}{\text{std}\left\{r^{(1)},r^{(2)},\ldots,r^{(G)}\right\}}$.
To guide the model toward producing responses with higher advantages, we update the policy $\pi_{\theta}$ by maximizing the following objective:
@@
Reward Design. Unlike discrete actions or text tokens, visual outputs are sparse, high-dimensional, and not easily decomposable into interpretable units. In our visual planning framework, the challenge is even more specific: whether the generated visual state can correctly reflect the intended planning action. Therefore, the reward design focus on the progress toward the goal while validating the actions with constraints. To interpret the intended action that connects the current state $v_{i}$ to a generated candidate state $\hat{v}_{i+t}^{(k)}$, we define a state-action parsing function $\mathcal{P}:\mathcal{V}\times\mathcal{V}\rightarrow\mathcal{A}\cup\mathcal{E}$, where $\mathcal{A}$ denotes the set of valid actions, and $\mathcal{E}$ is the set of invalid transitions, such as a violation of physical constraints of the environment. Formally,
$\mathcal{P}(v_{i},\hat{v}_{i+1}^{(k)})=\begin{cases}a_{i}^{(k)},&\text{if }a_{i}^{(k)}\in\mathcal{A},\\ e_{i}^{(k)},&\text{if }e_{i}^{(k)}\in\mathcal{E}.\end{cases}$ (4)
It helps to interpret model behaviors from pixel data to intended action through either standalone segmentation components [41] or rule-based scripts. Once having the intended actions, to systematically evaluate action effectiveness, we introduce the progress map $D(v)\in\mathbb{N}$ that estimates the remaining steps or effort required to reach the goal from each visual state. By comparing the agent’s current and resulting state against the progress map, we partition $\mathcal{A}\cup\mathcal{E}$ into three disjoint subsets:
$\mathcal{A}_{\mathrm{opt}}=\bigl{\{}a\in\mathcal{A}:D(\hat{v}_{i+1}^{(k)})<D(v_{i})\bigr{\}},\quad\mathcal{A}_{\mathrm{nopt}}=\bigl{\{}a\in\mathcal{A}:D(\hat{v}_{i+1}^{(k)})\geq D(v_{i})\bigr{\}},\quad\mathcal{E}_{\mathrm{inv}}=\mathcal{E}.$
We then propose the progress reward function $r(v_{i},\hat{v}_{i+1}^{(k)})$ as:
$\underbrace{\alpha_{\text{opt}}\cdot\mathbb{I}\!\bigl{[}\mathcal{P}(v_{i},\hat{v}_{i+1}^{(k)})\in\mathcal{A}_{\mathrm{opt}}\bigr{]}}_{\text{optimal}}+\;\underbrace{\alpha_{\text{nopt}}\cdot\mathbb{I}\!\bigl{[}\mathcal{P}(v_{i},\hat{v}_{i+1}^{(k)})\in\mathcal{A}_{\mathrm{nopt}}\bigr{]}}_{\text{non-optimal}}+\;\underbrace{\alpha_{\text{inv}}\cdot\mathbb{I}\!\bigl{[}\mathcal{P}(v_{i},\hat{v}_{i+1}^{(k)})\in\mathcal{E}_{\mathrm{inv}}\bigr{]}}_{\text{invalid}},$ (5)
@@ 3 Experiments and Results
Tasks
To evaluate our proposed visual planning paradigm, we select representative tasks where planning can be expressed and executed entirely in the visual modality. We focus on tasks where state transitions are visually observable, distinguishing them from language-centric tasks like code generation [27] or traditional visual question answering. This design allows us to analyze planning behavior without relying on textual rationales or symbolic outputs. To compare visual planning with language-based reasoning, we experiment with 3 visual navigation environments: FrozenLake [53], Maze [23], and MiniBehavior [25]. All of them can be solved in both modalities, which enables a direct comparison between visual planning and language reasoning strategies.
@@
[…]
Models
To explore visual planning without any language influence as confounders and enables a clean investigation, we select models trained exclusively on visual data without any exposure to textual data during pretraining. For our methods (VPFT and VPRL), we use the Large Vision Model (LVM-3B) [4] as the backbone, which is only trained on image sequences and videos. For RL training, we design and provide the detailed implementation of rule-based state-action parsing function $\mathcal{P}$ and progress map $D(v)$ in Appendix B.3.
We also include textual planning baselines for parallel comparison, where planning is formulated through language, typically as a textual sequence of actions. Specifically, we evaluate Qwen 2.5-VL-Instruct [3], matched in size to LVM-3B, on both inference-only (Direct [^1] Direct denotes answer prediction without being instructed to conduct intermediate reasoning. and CoT) and post-training settings (SFT) as baselines. We further evaluate closed-source models, including Gemini 2.0 Flash [26] (gemini-2.0-flash-002) and advanced thinking model Gemini 2.5 Pro [13] (gemini-2.5-pro-preview-03-25) as a reference from state-of-the-art multimodal reasoning. Full training details and hyperparameters for all models are provided in Appendix B.4.
Evaluation Metrics
We adopt two complementary evaluation metrics for the selected tasks:
  • Exact Match (EM) is defined as $\mathrm{EM}_{i}=\prod_{j=1}^{n}\mathbb{I}(\hat{v}_{j}=v_{j})$. This metric measures whether the model successfully generates the complete and correct planning trajectory that aligns with the shortest optimal valid path. One step of deviation from the optimal solution is considered incorrect.
  • Progress Rate (PR) is defined as $\mathrm{PR}_{i}=\frac{1}{n}\sum_{j=1}^{n}\left[\prod_{k=1}^{j}\mathbb{I}(\hat{v}_{k}=v_{k})\right]$. $\mathrm{PR}$ measures the ratio of the number of consecutively correct steps (valid forward moves) from the start to the number of steps in the optimal path. This provides a softer signal than Exact Match, capturing the model’s ability to make meaningful progress towards a full solution.
@@
Visual Planning Surpasses Textual Planning. Table 1 shows that visual planners (VPFT and VPRL) achieve the highest scores on all tasks, outperforming all language-reasoning baselines. With identical supervised training method via fine-tuning, VPFT exceeds language-based SFT by an average of over 22% in Exact Match (EM), with VPRL further widening the gap. A similar trend is observed in Progress Rate (PR) as well. This highlights the advantages of the visual planning paradigm in visual-centric tasks, where language-driven approaches may be less aligned with task structure. Inference-only models, whether large closed-source systems or smaller open-source MLLMs, struggle with these planning tasks without task-specific tuning. Even the advanced thinking model Gemini 2.5 Pro achieves EM and PR almost below 50% on the more complex Maze and MiniBehavior tasks, underscoring the challenges these tasks pose for current models despite being intuitive for humans.
Gains from Reinforcement Learning. The two-stage reinforcement learning approach (VPRL) yields the highest overall performance, surpassing all system variants. After Stage 2, the model achieves near-perfect planning on the simpler FrozenLake task (91.6% EM, 93.2% PR) and maintains strong performance on Maze and MiniBehavior tasks. This marks a substantial improvement of more than 20% across all tasks over the supervised baseline VPFT. As expected, Stage 1 of our RL training, which enforces output format without teaching planning behavior, yields near-random performance (e.g., 11% EM on FrozenLake, see Table 8 in Appendix C.5). After the full Stage 2 optimization with our reward scheme, the planner achieve its best performance. This gain highlights a key advantage of RL over SFT. VPRL allows the model to freely explore diverse actions and learn from their outcome, while VPFT relies on imitation and tends to fit the training distribution. By encouraging exploitation through reward-driven updates, VPRL learns to capture underlying rules and patterns, leading to stronger planning performance.
@@
Robustness with Scaling Complexity. The advantage of RL also holds when we study the performance of different methods with respect to task difficulties, where a larger grid usually relates to higher difficulties. In Figure 6, as the grid size increases from $3\times 3$ to $6\times 6$ in the FrozenLake environment, Gemini 2.5 Pro’s EM score drops sharply from 98.0% to 38.8%. In comparison, our visual planners not only maintain higher accuracy at all grid sizes but also exhibit a much flatter performance curve. Similarly, VPRL demonstrates even greater stability than VPFT, with EM remaining at 97.6% on $3\times 3$ grids and still achieving 82.4% on $6\times 6$, indicating strong robustness. We observe similar trends in other tasks; see Appendix C.2 for other tasks.

### v2

@@ Abstract
Recent advancements in Large Language Models (LLMs) and their multimodal extensions (MLLMs) have substantially enhanced machine reasoning across diverse tasks. However, these models predominantly rely on pure text as the medium for both expressing and structuring reasoning, even when visual information is present. In this work, we argue that language may not always be the most natural or effective modality for reasoning, particularly in tasks involving spatial and geometrical information. Motivated by this, we propose a new paradigm, Visual Planning, which enables planning through purely visual representations for these “vision-first” tasks, as a supplementary channel to language-based reasoning. In this paradigm, planning is executed via sequences of images that encode step-by-step inference in the visual domain, akin to how humans sketch or visualize future actions. We introduce a novel reinforcement learning framework, Visual Planning via Reinforcement Learning (VPRL), empowered by GRPO for post-training large vision models, leading to substantial improvements in planning in a selection of representative visual navigation tasks, FrozenLake, Maze, and MiniBehavior. Our visual planning paradigm outperforms all other planning variants that conduct reasoning in the text-only space. Our results establish Visual Planning as a viable and promising supplement to language-based reasoning, opening new avenues for tasks that benefit from intuitive, image-based inference. Code is available at: https://github.com/yix8/VisualPlanning.
@@ 1 Introduction
Building on this observation, we argue that performing multimodal reasoning only in the text pathway may not always offer the most intuitive or effective strategy, particularly for tasks that depend heavily on visual information and/or are ‘vision-first’ by design. Indeed, recent results from multimodal benchmarks [47, 33, 8, 9] offer growing evidence that purely language-based reasoning falls short in certain domains, particularly those involving spatial, geometric, or physical dynamics [60]. Such reliance on grounding visual information into text before reasoning introduces a modality gap that hinders the model’s ability to capture visual features and state transitions. This highlights a potential shortcoming of current MLLMs: while they process image inputs, they do not naturally “think” in images. For instance, tasks such as planning a route through a maze, designing the layout of a room, or predicting the next state of a mechanical system are often better served by visual representations, as verbal descriptions may be less effective and struggle to accurately capture complex spatial reasoning relationships. These examples suggest a broader question, which we aim to tackle in this work: can models directly plan in non-verbal modalities, such as images, without being mediated by text?
@@
We validate the feasibility of our paradigms on grid-based navigation as a representative of spatial planning tasks, including Maze [26], FrozenLake [57], and MiniBehavior [28], where one agent is requested to navigate to a target location successfully without violating environment constraints. Our experiments reveal that the visual planning paradigm substantially surpasses the traditional textual reasoning method by supervised fine-tuning (SFT), achieving 27% higher average exact-match rate. In addition to better performance, our novel method VPRL exhibits stronger generalization to out-of-distribution scenarios than the SFT method in the visual planning paradigm (VPFT). To the best of our knowledge, we are the first to apply RL to image generation in the context of planning, with contributions as follows:
@@ 2.1 The Visual Planning Paradigm
The majority of prior visual reasoning benchmarks [15, 1, 59] can be and is typically tackled by grounding the visual information in the textual domain [19, 43, 61], followed by a few steps of textual reasoning. However, once the visual content is mapped to text (e.g., object names, attributes, or relations), the problem gets reduced to a language reasoning task, where the reasoning is carried out by the language model, even without reflecting any information from the visual modality.
Our visual planning paradigm is fundamentally different. It performs planning purely within the visual modality as a holistic process, where the actions are not explicitly predicted but instead implicitly represented by transitions between visual states. We formally define visual planning as a process of generating a sequence of intermediate images $\hat{\mathcal{T}}=(\hat{v}_{1},\ldots,\hat{v}_{n})$, where each $\hat{v}_{i}$ represents a visual state that together constitute a visual planing trajectory, given the input image $v_{0}$. Specifically, let $\pi_{\theta}$ denote a generative vision model parameterized by $\theta$. The visual planning trajectory $\hat{\mathcal{T}}$ is generated autoregressively, where each intermediate visual state $\hat{v}_{i}$ is sampled conditioned on the initial state and previously generated states:
$\hat{v}_{i}\sim\pi_{\theta}(v_{i}|v_{0},\hat{v}_{1},...,\hat{v}_{i-1}).$ (1)
@@ 2.2 Reinforcement Learning for Large Vision Models
Reinforcement learning (RL) has shown notable advantages in improving the generalization of autoregressive models by optimizing with sequence-level rewards beyond token-level supervision signals [12]. In autoregressive image generation, an image is represented as a sequence of visual tokens. Inspired by the success of RL in language reasoning [17], we introduce an RL-based training framework for visual planning empowered by Group Relative Policy Optimization (GRPO) [48]. It leverages the transitions between visual states to compute the reward signals while verifying the constraints from the environments. To enforce the policy model that generates valid actions with diverse exploration during the RL process, we then propose a novel two-stage reinforcement learning framework for visual planning. In Stage 1, we first apply supervised learning to initialize the policy model with random trajectories. Model’s visual planning is then optimized by the RL training in Stage 2.
Stage 1: Policy Initialization. In this stage, we initialize the model $\pi_{\theta}$ by training it on random trajectories obtained by random walks in the environment. The goal here is to generate valid sequences of visual states and retain exploration capability in a ‘simulated’ environment. For training, each trajectory $\mathcal{T}$ consists of a sequence of visual states $(v_{0},\ldots,v_{n})$. From each trajectory, we extract n1n-1 image pairs of the form $(v_{\leq i},v_{i+1})$, where $v_{\leq i}$ represents the prefix sequence $(v_{0},\ldots,v_{i})$. Given an input prefix $v_{\leq i}$, to prevent overfitting to the specific transition and encourage stochasticity, we randomly sample one candidate state $\tilde{v}_{i+1}$ from all possible valid next states as the supervision target, and minimize the following loss function of visual planning via fine-tuning (VPFT):
$\mathcal{L}_{\text{VPFT}}(\theta)=-\mathbb{E}_{(v_{\leq i},\,\tilde{v}_{i+1})}\Bigl[\log\pi_{\theta}\!\bigl(\tilde{v}_{i+1}\,\big|\,v_{\leq i}\bigr)\Bigr].$ (2)
@@
Stage 2: Reinforcement Learning for Visual Planning. Building on Stage 1, where the model is initialized with random trajectories, it acquires the effective exploration capability. This property is essential for RL, as it ensures coverage over all possible transitions and prevents collapse to suboptimal behaviors. Stage 2 then leverages this ability to simulate the outcomes of potential actions by generating the next visual state and guiding the model to effectively do the planning. During this stage, the RL algorithm provides feedback and rewards based on the correctness of the simulated actions, gradually enabling the model to learn effective visual planning.
Specifically, given an input prefix $v_{\leq i}$, the behavior model $\pi_{\theta_{\text{old}}}$ samples a group of $G$ candidate responses $\{\hat{v}_{i+1}^{(1)},\ldots,\hat{v}_{i+1}^{(G)}\}$. The candidate response is then scored using a composite reward function $r(v_{i},\hat{v}_{i+1}^{(k)})$, which quantifies whether the generated visual state represents meaningful progress toward the goal state. The reward design and implementations are described in detail in the next paragraph.
Instead of relying on a learned critic to estimate value functions which may introduce additional sources of uncertainty and complexity, GRPO provides more computationally efficient and interpretable training signals by computing relative advantages through comparisons within the group. In this case, the relative advantage of each candidate is $A^{(k)}=\frac{r^{(k)}-\text{mean}\left\{r^{(1)},r^{(2)},\ldots,r^{(G)}\right\}}{\text{std}\left\{r^{(1)},r^{(2)},\ldots,r^{(G)}\right\}}$. To guide the model toward producing responses with higher advantages, we update the policy $\pi_{\theta}$ by maximizing the following objective:
@@
Reward Design. Unlike discrete actions or text tokens, visual outputs are sparse, high-dimensional, and not easily decomposable into interpretable units. In our visual planning framework, the challenge is even more specific: whether the generated visual state can correctly reflect the intended planning action. Consequently, our reward design emphasizes both adherence to environment constraints (validity of state transitions) and progress toward the goal.
Formally, let $\mathcal{A}$ denote the set of valid actions and $\mathcal{E}$ the set of invalid ones (e.g., violations of physical constraints or hallucinated new entities in the environment). To interpret and evaluate the intended action that connects the current state $v_{i}$ to a generated candidate state $\hat{v}_{i+t}^{(k)}$, we introduce 1) the dynamics interpreter $\mathcal{D}:\mathcal{V}\times\mathcal{V}\rightarrow\mathcal{A}\cup\mathcal{E}$ to parse the transition and 2) the progress estimator $P:\mathcal{V}\rightarrow\mathbb{N}$ to quantify the progress.
The dynamics interpreter $\mathcal{D}$ evaluates the transitions $a\in\mathcal{A}\cup\mathcal{E}$ for validity, which, by implementation, can be a dynamics model [44] or a rule-based system to elicit actions from state pairs, or a neural model as holistic validator that judges transitions without explicitly inferring actions. The progress estimator $P(v)$ quantifies progress by estimating the remaining steps or effort required to reach the goal from each visual state. By comparing the agent’s current and predicted state, we partition the generated candidate states $\mathcal{A}\cup\mathcal{E}$ into three disjoint subsets:
$\mathcal{A}_{\mathrm{opt}}=\bigl\{a\in\mathcal{A}:P(\hat{v}_{i+1}^{(k)})<P(v_{i})\bigr\},\quad\mathcal{A}_{\mathrm{nopt}}=\bigl\{a\in\mathcal{A}:P(\hat{v}_{i+1}^{(k)})\geq P(v_{i})\bigr\},\quad\mathcal{E}_{\mathrm{inv}}=\mathcal{E}.$
Here, $\mathcal{A}_{\mathrm{opt}}$ corresponds to optimal actions that reduce the distance to the goal, $\mathcal{A}_{\mathrm{nopt}}$ captures non-optimal but still valid actions, and $\mathcal{E}_{\mathrm{inv}}$ denotes invalid ones determined by the dynamics interpreter.
Based on this partition, we define the progress reward function $r(v_{i},\hat{v}_{i+1}^{(k)})$:
$\underbrace{\alpha_{\text{opt}}\cdot\mathbb{I}\!\bigl[\mathcal{D}(v_{i},\hat{v}_{i+1}^{(k)})\in\mathcal{A}_{\mathrm{opt}}\bigr]}_{\text{optimal}}+\;\underbrace{\alpha_{\text{nopt}}\cdot\mathbb{I}\!\bigl[\mathcal{D}(v_{i},\hat{v}_{i+1}^{(k)})\in\mathcal{A}_{\mathrm{nopt}}\bigr]}_{\text{non-optimal}}+\;\underbrace{\alpha_{\text{inv}}\cdot\mathbb{I}\!\bigl[\mathcal{D}(v_{i},\hat{v}_{i+1}^{(k)})\in\mathcal{E}_{\mathrm{inv}}\bigr]}_{\text{invalid}},$ (4)
@@ 3 Experiments and Results
Tasks. To evaluate our proposed visual planning paradigm, we select representative tasks where planning can be expressed and executed entirely in the visual modality. We focus on tasks where state transitions are visually observable, distinguishing them from language-centric tasks like code generation [30] or traditional visual question answering. This design allows us to analyze planning behavior without relying on textual rationales or symbolic outputs. To compare visual planning with language-based reasoning, we experiment with 3 visual navigation environments: FrozenLake [57], Maze [26], and MiniBehavior [28]. All of them can be solved in both modalities, which enables a direct parallel comparison of pros and cons between visual planning and language reasoning strategies.
@@
Table 1: Performance of the closed- and open-source models on FrozenLake, Maze, and MiniBehavior. VPRL performs consistently the best (bold) across all tasks. denotes the post-trained model. ˜ represents texts and  represents images. The last column Avg. reports the average performance across three tasks.
Model Input Output FrozenLake Maze MiniBehavior Avg. EM (%) PR (%) EM (%) PR (%) EM (%) PR (%) EM (%) PR (%) Closed-Source Model Gemini 2.0 Flash   xxx- Direct ˜+ ˜ 21.2 47.6 8.3 31.4 0.7 29.8 10.1 36.3   xxx- CoT ˜+ ˜ 27.6 52.5 6.9 29.8 4.0 31.2 12.8 37.8 Gemini 2.5 Pro (think) ˜+ ˜ 72.0 85.0 21.5 35.5 37.6 59.9 43.7 60.1 Open-Source Model Qwen 2.5-VL-Instruct-7B   xxx- Direct ˜+ ˜ 1.2 15.0 0.6 14.5 0.3 9.8 0.7 13.1   xxx- CoT ˜+ ˜ 8.2 29.1 2.3 15.2 0.5 14.7 3.7 19.7   xxx- SFT ˜+ ˜ 68.6 84.4 60.9 70.3 31.3 56.1 53.6 69.9 LVM-7B   xxx- VPFT (ours) 75.4 79.5 59.0 64.0 33.8 52.2 56.1 65.2   xxx- VPRL (ours) 91.6 93.2 74.5 77.6 75.8 83.8 80.6 84.9
Models. To explore visual planning without any language influence as confounders and enables a clean investigation, we select models trained exclusively on visual data without any exposure to textual data during pretraining. For visual planning, we use the Large Vision Model (LVM-7B) [4] as the backbone, which is only trained on image sequences and videos. We train the model with 1) supervised fine-tuning over golden planning trajectory (VPFT) and 2) two-stage reinforcement learning (VPRL), resulting in two system variants with visual planning. For RL training, we start with a rule-based parsing function as the dynamics interpreter to parse the image pairs to actions, and a heuristic progress estimator, with details enclosed in Appendix B.3.
For baselines, to facilitate parallel comparison for language-based planning, we adopt Qwen 2.5-VL-Instruct [3], on both inference-only (Direct [^1] Direct denotes answer prediction without being instructed to conduct intermediate reasoning. and CoT) and post-training settings (SFT and RL), trained on the same data as the visual planner. We further evaluate multimodal reasoning performance of proprietary models with Gemini 2.0 Flash [29] and advanced thinking model Gemini 2.5 Pro [14]. Full training details, model versions, and hyperparameters are provided in Appendix B.4.
Evaluation Metrics. We adopt two complementary evaluation metrics for the selected tasks. Let $\mathcal{O}=\{\mathcal{T}^{(1)},\mathcal{T}^{(2)},\dots,\mathcal{T}^{(M)}\}$ denote the set of all shortest optimal trajectories of length $n$, where each trajectory is $\mathcal{T}^{(m)}=(v^{(m)}_{1},\dots,v^{(m)}_{n})$, and let $\hat{\mathcal{T}}=(\hat{v}_{1},\dots,\hat{v}_{n})$ denote the predicted trajectory.
  • Exact Match (EM) is defined as $\mathrm{EM}=\max_{m\in\{1,\dots,M\}}\prod_{j=1}^{n}\mathbb{I}(\hat{v}_{j}=v^{(m)}_{j})$, evaluating whether $\hat{\mathcal{T}}$ coincides with any $\mathcal{T}^{(m)}\in\mathcal{O}$. EM requires the entire trajectory to be valid and of minimal length, and accepts all optimal solutions rather than a single reference.
  • Progress Rate (PR) is defined as $\mathrm{PR}=\max_{m\in\{1,\dots,M\}}\frac{1}{n}\sum_{j=1}^{n}\left[\prod_{k=1}^{j}\mathbb{I}(\hat{v}_{k}=v^{(m)}_{k})\right$], measuring the ratio of consecutive correct steps (valid forward moves) from the start that align with at least one optimal trajectory. PR thus provides a softer signal than EM, capturing the model’s ability to make meaningful progress towards a full solution.
@@
Table 2: Performance of text-based planning variants on FrozenLake. See Table 7 in Appendix C.2 for the full results.
Model EM (%) PR (%)
Qwen 2.5-VL-Instruct-7B
- SFT
  xxx- Direct 68.6 84.4
  xxx- w/ Coordinates 74.4 82.7
  xxx- w/ ASCII 73.1 83.4
- GRPO
  xxx- w/ VPRL progress reward 54.4 69.9
  xxx- w/ PR metric reward 60.1 74.3
Textual planning falls short in both proprietary models and open-sourced tuning baselines. Table 1 shows that proprietary models yield average EM below 50% and PR only marginally above 50% at best, underscoring the challenges these tasks pose for current models despite being intuitive for humans. On the other hand, while task-specific training provides partial improvement, the overall performance of fine-tuned textual planners remains unsatisfactory, through either directly generating planned actions (SFT in Table 1) or first captioning the image with different textual representations and then generating answers (Table 2). We also observe that, unlike the notable gains of RL in the pure language domain [17], RL yields limited performance gains when applied to text-based planning with multimodal inputs. Table 2 shows that when using progress reward as in VPRL or directly using the Progress Rate metric as the outcome reward, none of the variants surpasses the SFT baseline. We attribute the bottleneck of language-based planning with SFT and RL to the modality gap, which leads to inaccuracies in grounding visual information into text and thereby constrains performance. Further discussion is provided in Section 4.
Visual planning achieves better performance than textual baselines via RL. While supervised fine-tuning (VPFT) achieves performance comparable to text-based SFT, it remains constrained by imitation and limited exposure to diverse trajectories. By contrast, our two-stage reinforcement learning framework (VPRL) substantially improves the planning capability, achieving the strongest overall results. After Stage 2 optimization, the model attains near-perfect accuracy on FrozenLake (91.6% EM, 93.2% PR) and maintains strong performance on more complex Maze and MiniBehavior tasks, outperforming VPFT by over 20% on average. As expected, the improvement is fully driven by outcome-based optimization in Stage 2, as Stage 1 alone yields near-random behavior (Table 10 in Appendix C.6). Unlike VPFT, which mainly fits the training distribution, VPRL enables exploration of diverse actions and learning from their consequences through reward-driven updates, thereby capturing underlying planning rules and achieving stronger performance.
@@
VPRL shows robustness with scaling complexity. The advantage of RL also holds when we study the performance of different methods with respect to task difficulties, where a larger grid usually relates to higher difficulties. In Figure 6, as the grid size increases from $3\times 3$ to $6\times 6$ in the FrozenLake environment, Gemini 2.5 Pro’s EM score drops sharply from 98.0% to 38.8%. In comparison, our visual planners not only maintain higher accuracy at all grid sizes but also exhibit a much flatter performance curve. Similarly, VPRL demonstrates even greater stability than VPFT, with EM remaining at 97.6% on $3\times 3$ grids and still achieving 82.4% on $6\times 6$, indicating strong robustness. We observe similar trends in other tasks; see Appendix C.3 for other tasks.